Suppose the interest rate is 8% and a bond with annual coupon of $90 matures in one year time paying its face value of $1000. What is bond's current price?
1
Expert's answer
2020-06-11T10:54:46-0400

The price of a bond is calculated as:



"P = \sum\dfrac{C_t}{(1 + YTM)^t} + \dfrac{P}{(1 + r)^t}"

Therefore, the bond price is:



"P = \dfrac{90}{(1 + 0.08)^1} + \dfrac{1000}{(1 + 0.08)^1}\\[0.3cm]
\color{red}{P = \$1,009.26}"